OpenStudy (anonymous):

timber increase by 2% each year. if a consignment of timber is currently worth $100 000 a) express the value of timber, P dollars, as a function of time, t, where t is the no. of years from now.

OpenStudy (anonymous):

Pwease help!

OpenStudy (anonymous):

P = 100,000 + 2% 100,000 * t = 100,000 ( 1 + 1,000 t)

OpenStudy (anonymous):

in my text book the answer is P = 100,000(1.02)*t

OpenStudy (anonymous):

@chlorophyll

OpenStudy (anonymous):

OOps, calculating mistakes :") Yep, 100,000 factor out, so inside is 1 + .02 = 1.02

OpenStudy (anonymous):

P = 100,000(1.02)*t The idea is increment over time!

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yup thankyou!!!

OpenStudy (anonymous):

I'm glad that you ask :)